### v1

# Stylos: Multi-View 3D Stylization with Single-Forward Gaussian Splatting

Abstract
We present Stylos, a single-forward 3D Gaussian framework for 3D style transfer that operates on unposed content, from a single image to a multi-view collection, conditioned on a separate reference style image. Stylos synthesizes a stylized 3D Gaussian scene without per-scene optimization or precomputed poses, achieving geometry-aware, view-consistent stylization that generalizes to unseen categories, scenes, and styles. At its core, Stylos adopts a Transformer backbone with two pathways: geometry predictions retain self-attention to preserve geometric fidelity, while style is injected via global cross-attention to enforce visual consistency across views. With the addition of a voxel-based 3D style loss that aligns aggregated scene features to style statistics, Stylos enforces view-consistent stylization while preserving geometry. Experiments across multiple datasets demonstrate that Stylos delivers high-quality zero-shot stylization, highlighting the effectiveness of global style–content coupling, the proposed 3D style loss, and the scalability of our framework from single view to large-scale multi-view settings. Our codes will be available at https://github.com/HanzhouLiu/StylOS.
1 Introduction
Image guided 3D stylization aims to preserve scene geometry and cross-view consistency while transferring the reference style. With the rise of immersive content, augmented and virtual reality, demand for this capability is growing. Nevertheless, achieving reliable 3D stylization remains challenging and continues to attract significant research attention.
Early attempts rely on explicit 3D geometry such as meshes or point clouds, but their stylization quality is often limited by geometric reconstruction accuracy, resulting in noticeable artifacts in complex scenes (Kato et al., 2018; Cao et al., 2020). The introduction of implicit radiance fields, most notably NeRF (Mildenhall et al., 2020), enables higher-fidelity rendering, and subsequent works extended it to artistic stylization (Zhang et al., 2022; Huang et al., 2022; Liu et al., 2023), but typically require costly per-scene optimization, constraining their generalization to unseen scenes. More recently, 3D Gaussian Splatting (3DGS) has emerged as a promising explicit representation that combines high reconstruction quality with real-time rendering efficiency (Kerbl et al., 2023). Stylization approaches built on 3DGS achieve efficient multi-view consistency (Liu et al., 2024; Galerne et al., 2025), yet still struggle to generalize beyond scene-specific training.
In contrast, we introduce Stylos (meaning pens in French), a single-forward framework for 3D style transfer that eliminates the need for per-scene optimization and precomputed camera parameters, and generalizes effectively to novel categories, scenes, and styles. Stylos employs a shared Transformer backbone with two pathways: content and style images are projected into a shared feature space, where content retains self-attention for geometric reasoning, and style is injected via global cross-attention to condition aesthetics. Notably, the global cross-attention also benefits the preservation of geometric details. Geometry-related attributes, such as depth, camera intrinsics, and extrinsics, are derived from backbone features, whereas style conditioning guides the prediction of color coefficients. These outputs are estimated through prediction heads that serve as the interface between feature space and the final Gaussian representation. Prior works in 2D style transfer typically rely on distribution alignment (Li et al., 2017), which operate on image statistics but lack explicit multi-view or 3D consistency. To address this, we explore alternative objectives and propose a voxel-based 3D style loss that aligns aggregated scene features with style statistics, providing stronger view-consistent stylization while preserving geometric fidelity.
We evaluate Stylos across different scenarios, including category-level transfer and cross-scene generalization. Our assessment spans challenging real-world benchmarks, where Stylos produces stylized renderings with high visual fidelity and consistent geometry, demonstrating robustness even in previously unseen environments. Our main contributions are threefold:
  • We propose a shared-backbone design with two pathways: geometry predictions retain self-attention for geometric reasoning, while style is injected through global cross-attention to condition aesthetics.
  • We introduce a voxel-based 3D style loss that aligns aggregated scene features with style statistics, enforcing cross-view coherence and geometry-aware stylization.
  • We develop Stylos, a single-forward-pass pipeline for 3D style transfer from unposed inputs, scaling from a single to hundreds of views with a single style image, and achieving zero-shot generalization to unseen categories, scenes, and styles.
2 Related Work
2.1 Single-Forward 3D Reconstruction
Early CNN models (Ummenhofer et al., 2017; Zhou et al., 2018; Teed & Deng, 2020) demonstrate the feasibility of end-to-end 3D reconstruction. While the Transformer-based DUSt3R (Wang et al., 2024) has shown impressive performance, yet remains constrained by pairwise inputs and reliance on traditional optimization for pose estimation. VGGT (Wang et al., 2025a) advances this line by jointly predicting camera parameters, depth, point maps, and tracks from one to hundreds of views in a single forward pass, while AnySplat (Jiang et al., 2025) complements it with a feed-forward Gaussian splatting pipeline trained with VGGT priors to directly predict Gaussian parameters along with depth and camera parameters from uncalibrated images.
2.2 3D Style Transfer and Losses
Explicit-representation approaches to 3D stylization (Kato et al., 2018; Cao et al., 2020; Guo et al., 2021; Huang et al., 2021; Höllein et al., 2022; Mu et al., 2022) have shown promising results, but their reliance on accurate geometry limits applicability in complex real-world scenes. Implicit neural representations like NeRF (Mildenhall et al., 2020) improve geometric consistency, inspiring a range of stylization methods based on stylized view supervision or auxiliary networks (Nguyen-Phuoc et al., 2022; Zhang et al., 2022; Chiang et al., 2022; Huang et al., 2022). StyleRF (Liu et al., 2023) further enhances fidelity and enables zero-shot transfer by applying style transformations in radiance-field feature space.
More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as an efficient alternative to NeRF. StyleGaussian (Liu et al., 2024) extends 3DGS with efficient feature rendering to enable real-time stylization, while subsequent works refine 3DGS for specialized settings and efficiency improvements (Mei et al., 2024; Zhang et al., 2024; Yu et al., 2024; Galerne et al., 2025). In parallel, Styl3R (Wang et al., 2025b) disentangles geometry and style using separate encoders, but is restricted to 2–8 input views. In contrast, Stylos follows the shared-backbone paradigm of AdaIN (Huang & Belongie, 2017), leveraging VGGT for unified feature extraction, and scales from a single to hundreds of input views.
Underlying these architectures, style transfer objectives play a crucial role. Classical style transfer relies on Gram-matrix correlations (Gatys et al., 2016), while channel-wise statistics such as mean and variance (Li et al., 2017) offer efficient alternatives, forming the basis of AdaIN (Huang & Belongie, 2017). Extensions to video and multi-view stylization (Ruder et al., 2016; Gupta et al., 2017; Nguyen-Phuoc et al., 2022; Zhang et al., 2022) address temporal and cross-view consistency, and CLIP-based losses (Radford et al., 2021) introduce semantic alignment with text or image prompts. However, most objectives remain inherently 2D and cannot guarantee consistency across complex 3D scenes. To this end, we extend statistics-based style losses into voxel space, introducing a 3D-aware objective that aligns feature distributions after multi-view fusion. This formulation complements perceptual and CLIP-based losses, enforcing geometry-aware and view-consistent stylization.
Figure 1: Architecture overview. Given multi-view content inputs and a style reference, Stylos enables instant 3D stylization without scene-specific training or post-optimization. A key component is the 3D style loss, matching voxelized 3D features with 2D style statistics.
3 Method
We propose Stylos, a transformer-based framework for stylized 3D scene reconstruction. Given a style reference image and one or more content views (also referred to as context views), Stylos predicts a set of stylized 3D Gaussian primitives together with camera parameters, enabling faithful reconstruction of the observed scene while transferring the desired style. We first formulate the problem in Sec. 3.1, then present the network architecture in Sec. 3.2, and finally detail our training strategy in Sec. 3.3 and the used style loss in Sec. 3.4.
3.1 Problem Formulation
We aim to disentangle geometry and style within a unified 3D scene representation. The input to Stylos consists of a set of $N$ context views $\{I_{i}\}_{i=1}^{N}$ of a scene, with $N\geq 1$ and $I_{i}\in\mathbb{R}^{H\times W\times 3}$, together with a single style reference image $S\in\mathbb{R}^{H\times W\times 3}$. The context views provide geometric cues, while the style image specifies the desired aesthetic.
Formally, Stylos defines a conditional mapping,
$f_{\theta}:(\{I_{i}\}_{i=1}^{N},\,S)\;\mapsto\;\big(G,\,\{g_{i}\}_{i=1}^{N}\big),$
where the scene representation $G=\{(p_{m},\,c_{m})\}_{m=1}^{M}$ is parameterized by $M$ anisotropic 3D Gaussians. Each Gaussian has geometry attributes $p_{m}=(\mu_{m},\,\alpha,\,r_{m},\,s_{m})$, comprising 3D position $\mu_{m}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}_{+}$, orientation quaternion $r_{m}\in\mathbb{R}^{4}$, and anisotropic scale $s_{m}\in\mathbb{R}^{3}$, as well as a style-dependent color embedding $c_{m}\in\mathbb{R}^{3\times(k+1)^{2}}$ represented with spherical harmonics of degree $k$. In addition, Stylos jointly predicts camera parameters $\{g_{i}\in\mathbb{R}^{9}\}_{i=1}^{N}$ for each input view, leveraging pose cues estimated by the VGGT backbone.
3.2 Network Architecture
Architecture Overview. At its core, Stylos employs a geometry backbone that alternates frame-wise and global attention to process multiple input views and infer geometry-related parameters such as positions, scales, orientations, and opacities. To enable stylization, we introduce a dedicated conditioning branch, the Style Aggregator, which fuses content and style features through cross-attention and predicts style-aware color embeddings for each Gaussian. In this way, geometry remains derived solely from the backbone, while style representation is conditioned on the style reference. We next detail the geometric backbone in Sec. 3.2.1, the style aggregation module in Sec. 3.2.2, and the prediction heads in Sec. 3.2.3.
3.2.1 Geometric Backbone
The geometric backbone follows the alternating-attention design of VGGT (Wang et al., 2025a), which interleaves frame and global self-attention layers to capture both intra-frame structure and cross-view consistency. This component is kept unchanged to retain strong geometric reasoning, serving as the foundation on which we integrate style information.
3.2.2 Style Aggregator
The Style Aggregator extends the geometric backbone by adapting its attention layers from self-attention to cross-attention. Both content views and the style image are embedded with a shared patch encoder, ensuring that their tokens lie in a common feature space. In this design, content tokens serve as queries while style tokens provide keys and values, so that style information conditions the token representations later used for predicting Gaussian colors. Geometry-related parameters, in contrast, are still predicted from the backbone pathway and remain unaffected by style conditioning.
Formally, let $T\in\mathbb{R}^{B\times N\times K\times C}$ denote the content tokens from $N$ views, each with $K$ patches and $C$ channels, and let $S\in\mathbb{R}^{B\times K\times C}$ represent the style tokens. Let $T_{i}$ be the token set of the $i$-th view, and let $\mathrm{Attn}(Q,KV)$ denote a standard cross-attention operation. Depending on how queries are formed, we distinguish three style–content coupling mechanisms,
(Frame-only): $T^{\text{frame}}_{i}=\mathrm{Attn}(T_{i},S),$ $Q=T_{i},\;\;K=V=S,$ (1)
(Global-only): $T^{\text{global}}=\mathrm{Attn}(\{T_{i}\}_{i=1}^{N},S),$ $Q=\{T_{i}\}_{i=1}^{N},\;\;K=V=S,$ (2)
(Hybrid): $T^{\text{hybrid}}=[\,\{T^{\text{frame}}_{i}\}_{i=1}^{N}\,\|\,T^{\text{global}}\,],$ concatenation along channel dim. (3)
Frame-only injects style independently per view, global-only enforces a shared conditioning across all views, and hybrid concatenates the frame-level and global-level outputs along the channel dimension, yielding $2\times C$-dimensional features. In our design, Stylos adopts the global-only configuration, implemented with CrossBlock layers where $\{T_{i}\}_{i=1}^{N}$ serve as queries and $S$ provides keys and values. The resulting features are passed to the prediction heads described in Sec. 3.2.3.
3.2.3 Prediction Heads
To connect the geometric backbone and Style Aggregator with the Gaussian scene, we introduce prediction heads that translate feature tokens into explicit parameters. These heads serve as modular interfaces, keeping geometry and style distinct in feature space while ensuring their integration in the final Gaussian representation. We next describe the individual heads in detail.
Geometry Head. Geometric backbone features are passed through a DPT-style regression head that outputs the Gaussian geometry parameters $p_{m}=(\mu_{m},s_{m},r_{m},\alpha_{m})$, i.e., position, scale, orientation, and opacity (as defined in Sec. 3.1). By relying on this established design, structural predictions are derived from backbone features alone, without direct influence from style conditioning.
Style Head. The outputs of the Style Aggregator are subsequently processed by a color head to predict the spherical-harmonic coefficients $c_{m}$ that define appearance. This pathway injects style information directly into Gaussian colors while leaving the geometry parameters $p_{m}$ unaffected, enabling the two factors to be recombined seamlessly at the Gaussian level.
Auxiliary Heads, Adapter, and Voxelization. We employ the existing VGGT camera head to estimate camera intrinsics and extrinsics, and a depth head to predict scene geometry cues, which are unprojected into 3D anchors for Gaussian placement (Kerbl et al., 2023; Ren et al., 2024). A Gaussian adapter then consolidates geometry and style outputs into a unified set of primitives $\{(p_{m},c_{m})\}_{m=1}^{M}$ for differentiable rendering. Finally, to reduce redundancy and balance density, we follow the voxelization step introduced by AnySplat (Jiang et al., 2025), nearby points are clustered within a discretized 3D grid and fused using confidence-aware weighting. This operation depends only on the unprojected 3D points and features, and is independent of camera parameters.
3.3 Training Strategy
We adopt a two-stage training strategy for structure-aware stylization.
Stage 1: Geometry Pretraining. We initialize the geometric backbone of Stylos with VGGT weights (Wang et al., 2025a) and train the network end-to-end to learn geometry and photometric appearance. To avoid trivial identity mapping and improve robustness to color variations, one input view is randomly selected and color-jittered as the style reference. A frozen VGGT teacher provides pose and depth supervision. The objective combines reconstruction and distillation,
$\mathcal{L}_{\text{stage1}}=\mathcal{L}_{\text{rec}}+\lambda_{\text{distill}}\,\mathcal{L}_{\text{distill}}.$
Stage 2: Stylization Fine-tuning. We freeze all geometry-related modules and only update the Style Aggregator and the color head. Following ArtFlow (An et al., 2021), we use feature-level style and content losses in VGG space, matching channel-wise statistics for style and feature activations for content. We further extend these objectives with a 3D voxel-space style loss for cross-view consistency, and add a CLIP-based loss for semantic alignment. A total variation (TV) regularizer is also included to suppress high-frequency artifacts and stabilize optimization. The total loss is,
$\mathcal{L}_{\text{stage2}}=\mathcal{L}_{\text{rec}}+\lambda_{\text{style}}\,\mathcal{L}_{\text{style}}^{3D}+\lambda_{\text{cnt}}\,\mathcal{L}_{\text{content}}+\lambda_{\text{clip}}\,\mathcal{L}_{\text{clip}}+\lambda_{\text{tv}}\,\mathcal{L}_{\text{TV}}.$
3.4 Style Losses
We denote by $F^{l}_{b,s}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the VGG feature map of the rendered (stylized) image for the $b$-th scene and $s$-th view at layer $l$, and by $S^{l}_{b}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the feature map of the style reference image. Building on the BN-statistics formulation (Li et al., 2017), which interprets channel-wise mean and variance as style descriptors, we progressively extend statistics matching beyond the standard 2D setting. This progression starts from independent image-level matching, moves to multi-view feature aggregation for enforcing cross-view consistency, and culminates in a voxel-space loss that directly constrains the fused 3D representation.
Image-Level Style Loss. The simplest baseline aligns each rendered frame independently with the style reference,
$\mathcal{L}^{\mathrm{img}}_{\mathrm{sty}}=\frac{1}{BS}\sum_{b=1}^{B}\sum_{s=1}^{S}\sum_{l}\alpha_{l}\Bigl(\|\mu(F^{l}_{b,s})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(F^{l}_{b,s})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (4)
This encourages per-frame stylization but does not enforce multi-view consistency.
Scene-Level Style Loss. To encourage consistent stylization across multiple views, we concatenate the per-view features $\{F^{l}_{b,s}\}_{s=1}^{S}$ along the spatial dimension to form $\widetilde{F}^{l}_{b}$, and compute statistics on this aggregated map,
$\mathcal{L}^{\mathrm{scn}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l}\alpha_{l}\Bigl(\|\mu(\widetilde{F}^{l}_{b})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(\widetilde{F}^{l}_{b})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (5)
This step enforces cross-view consistency but still operates in 2D feature space.
3D Style Loss. Finally, we fuse multi-view features into a voxel grid using differentiable unprojection, where features from different views are accumulated into spatial bins of a discretized 3D volume. Let $G^{l}_{b}$ denote the voxelized features for the $b$-th scene at layer $l$. We then compute style statistics directly in voxel space,
$\mathcal{L}^{\mathrm{3D}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l}\alpha_{l}\Bigl(\|\mu(G^{l}_{b})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(G^{l}_{b})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (6)
By operating on voxelized features, this loss explicitly encodes geometry and enforces style consistency across both views and the underlying 3D structure.
Figure 2: CO3D pizza scene comparing different style–content cross-attention strategies.
[…]
Strategies Skateboard Pizza Donut
[…]
21.12 0.6858 0.2821 19.78 0.5939 0.3326 21.39 0.7198 0.3264
20.93 0.6917 0.2912 19.72 0.5940 0.3405 21.40 0.7167 0.3340
21.68 0.7043 0.2684 20.57 0.6177 0.3110 22.09 0.7362 0.3125
4 Experiment
Datasets. We evaluate cross-category generalization on the CO3D (Reizenstein et al., 2021) dataset by training on 17 categories and testing on 3 held-out ones, and cross-scene generalization by training on the full DL3DV-10K (Ling et al., 2024) and testing on Tanks and Temples (Knapitsch et al., 2017b). Style supervision is provided by WikiArt (WikiArt, ), with 50 images reserved for unseen-style evaluation.
Baselines. We compare with two representative open-source baselines, StyleRF (Liu et al., 2023) and StyleGS (Liu et al., 2024), using their released per-scene trained weights against our single DL3DV-trained model. Concurrent to Stylos, the other instant stylization solution Styl3R (Wang et al., 2025b) was proposed, but its official codes are not yet available at the time of our submission. We plan to provide comparisons once they are released.
Evaluation Metrics. Our evaluation covers three aspects. (1) To assess geometry reconstruction in the style-free setting, we report PSNR, SSIM, and LPIPS (Zhang et al., 2018) between content views and predictions, conditioned on the original first frame of each scene as the style input. (2) To measure stylization consistency, following prior work (Chiang et al., 2022), we compute LPIPS and RMSE in both short-range (consecutive frames $t$ and t1t{-}1) and long-range (distant frames $t$ and t7t{-}7) settings. (3) To further evaluate stylization quality, we report ArtScore (Chen et al., 2024), a recent metric specifically designed for reference-free evaluation of artness in generated images.
4.1 Ablation Study
In this section, we analyze the impact of each design choice on the CO3D dataset. To ensure fairness, all variants are trained with identical configurations in each experimental round.
Style-Content Cross Attention. As reported in Table 1, restricting cross-attention to the frame level yields only moderate performance, with a PSNR of 20.9 dB on skateboard and 19.7 dB on pizza. The hybrid variant, which combines frame-level and global attention, provides observable improvement on donut with 21.4 dB, but brings limited gains on the other categories and overall underperforms the global-only setup. By contrast, applying cross-attention at the global level consistently achieves the best results, reaching a PSNR of 21.7 dB on skateboard, 20.6 dB on pizza, and 22.1 dB on donut. Qualitative comparisons in Fig. 2 highlight these differences on the pizza example. Frame-only attention produces a poorly defined crust with noticeable artifacts, the hybrid variant recovers part of the structure but leaves the box edges blurred and unstable, while global attention delivers the most faithful result, with clear toppings and a well-preserved crust boundary.
Figure 3: Comparison of style losses on unseen donut, skateboard, and pizza scenes from the CO3D dataset. Both scene and 3D style losses yield cleaner stylized textures compared to image-level matching, while the 3D loss further conveys a stronger sense of 3D geometry.
Table 2: Consistency comparison of different style loss designs on CO3D. The 3D loss achieves the strongest short-range and long-range consistency.
Style Loss Short-range Long-range
$downarrow LPIPS\downarrow RMSE\downarrow$
Image loss (Eq. 4) 0.046 0.035 0.147 0.135
Scene loss (Eq. 5) 0.048 0.037 0.147 0.133
3D loss (Eq. 6) 0.046 0.033 0.140 0.128
Image vs. Scene vs. 3D Style Losses. As shown in Table 2, the image-level loss achieves short-range LPIPS of 0.046 and RMSE of 0.035, but long-range consistency rises to 0.147 and 0.135. The scene-level loss slightly improves long-range RMSE to 0.133 but degrades short-range performance to 0.048 and 0.037, indicating that simple 2D aggregation reduces some long-range inconsistencies but compromises short-range accuracy. In contrast, the 3D loss applies statistics after voxel-space fusion, matching the best short-range LPIPS of 0.046, further lowering short-range RMSE to 0.033, and improving long-range scores to 0.140 and 0.128. These results indicate that enforcing style consistency in a geometry-aligned space yields sharper boundaries and more coherent textures across views. As further illustrated in Fig. 3, image-level supervision often fails to faithfully transfer style. For example, the donut lacks stylization and the wooden surface is not well synthesized, whereas scene- and 3D-level losses successfully apply the target style. Among them, the 3D style loss produces sharper boundaries and a stronger sense of 3D geometry, as observed in the skateboard and pizza examples, leading to the most coherent multi-view stylization.
From a Single View to Dozens of Views. While Stylos can process hundreds of content inputs, we qualitatively observe a gradual decrease in visual quality once the number of views per batch (denoted as “# views / batch”) exceeds 32. As shown in Fig. 4, very small batches (1 view) fail to synthesize parts of the scene such as the bench, while overly large batches (64 views) introduce edge artifacts on the tall building, potentially due to unstable Gaussian representations and the gap from training settings (no more than 24 views).
Figure 4: Effect of varying # views / batch on the Lighthouse scene from Tanks and Temples.
[…]
Method Train Truck M60 Garden
$downarrow LPIPS\downarrow RMSE\downarrow$ $downarrow LPIPS\downarrow RMSE\downarrow$
Short-range consistency
StyleRF 2023 0.075 0.065 0.060 0.062 0.039 0.042
StyleGS 2024 0.033 0.038 0.031 0.034 0.038 0.037 0.069 0.061
Stylos (ours) 0.030 0.026 0.028 0.021 0.035 0.024 0.047 0.044
Long-range consistency
StyleRF 2023 0.116 0.110 0.176 0.154 0.096 0.092
StyleGS 2024 0.067 0.072 0.086 0.077 0.091 0.091 0.177 0.141
Stylos (ours) 0.051 0.056 0.074 0.069 0.083 0.082 0.139 0.134
4.2 Comparison with State-of-the-Art
For StyleRF (Liu et al., 2023) and StyleGS (Liu et al., 2024), we directly evaluate with their released per-scene weights. In contrast, our Stylos is trained once on the full DL3DV dataset (Ling et al., 2024) and tested in a zero-shot manner, without ever seeing the images of these scenes. Quantitative Evaluation. According to Table 3, Stylos achieves consistently lower LPIPS and RMSE than StyleGS across both short- and long-range view pairs. This demonstrates that Stylos provides stronger cross-view consistency. Furthermore, Table 4 reports that Stylos achieves the highest ArtScore across nearly all four Tanks and Temples scenes (Knapitsch et al. (2017a)), outperforming both StyleRF and StyleGS.
[…]
Method Train Truck M60 Garden
StyleRF 2023 7.71 4.12 8.48
StyleGS 2024 0.78 5.76 8.63 9.38
Stylos (ours) 9.50 9.70 9.37 9.34
Figure 5: Visual comparison between Stylos and recent per-scene 3D stylization baselines.
Qualitative Evaluation. Fig. 5 shows that Stylos successfully transfers the target style while preserving scene geometry and delivering consistent textures. In contrast, StyleGS introduces noticeable geometric distortions and texture discontinues, while StyleRF tends to over-impose style patterns, negatively impacting the object structure. Overall, Stylos achieves the best trade-off between structural fidelity and faithful stylization.
5 Conclusion
In this work, we propose Stylos, a single-forward-pass framework for multi-view 3D stylization. By keeping geometry predictions on the self-attentive backbone path and conditioning color via global cross-attention, together with a voxel-space 3D style loss that aggregates multi-view features, Stylos achieves geometry-aware and view-consistent stylization. Our ablations show that the global-only cross-attention between style and content better preserves geometric details than alternative coupling strategies. Extensive experiments across category-level and large-scale scene datasets demonstrate the generalization ability of our approach, achieving aesthetically pleasing stylization with strong cross-view consistency. In the future, we aim to scale Stylos to support higher-resolution inputs while improving efficiency, paving the way for practical 3D content creation.
6 Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. The main paper describes all model architectures in Sec. 3. We release detailed pseudo-code for all loss functions in Appendix A, and will release the full implementation upon acceptance. However, to support reproducibility, we provide detailed pseudo codes of the proposed style loss functions (image-level, scene-level, and 3D-level) in Appendix A.1. All datasets used are publicly available, as described in Sec. 4. These resources should allow researchers to reproduce and extend our findings. In addition, we provide extensive qualitative results in Appendix A.3, and analyze efficiency trends with respect to the number of views per batch in Appendix A.2. Efficiency experiments are averaged over 100 iterations per setting.
Appendix A Appendix
A.1 Style Loss Pseudo Codes
We provide pseudo codes to the three variants for style losses introduced in the main paper, image-level, scene-level, and 3d style losses. This is for reproducibility purpose, and please follow the implementation in our codes which will be released in the near future.
Algorithm 1 Image-level AdaIN Style Loss
0: Content images $I_{c}\in\mathbb{R}^{B\times S\times C\times H\times W}$, Rendered images $I_{r}$, Style images $I_{s}$
0: Image-level style loss $\mathcal{L}_{\text{img}}$
1:for $i=1$ to $B$ do
2:  Extract style features $F_{s}\leftarrow\text{VGG}(I_{s}[i])$
3:  for $j=1$ to $S$ do
4:   Extract content features $F_{c}\leftarrow\text{VGG}(I_{c}[i,j])$
5:   Extract rendered features $F_{r}\leftarrow\text{VGG}(I_{r}[i,j])$
6:   Content loss:
$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
7:   Style loss initialization: $\mathcal{L}_{s}\leftarrow 0$
8:   for $k=1$ to 55 do
9:    $(\mu_{r},\sigma_{r})\leftarrow\text{MeanStd}(F_{r}^{(k)})$
10:    $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
11:    $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{r}-\mu_{s}\|_{2}^{2}+\|\sigma_{r}-\sigma_{s}\|_{2}^{2}$
12:   end for
13:  end for
14:end for
15:return $\mathcal{L}_{\text{img}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
Algorithm 2 Scene-level AdaIN Style Loss
0: Content images $I_{c}\in\mathbb{R}^{B\times S\times C\times H\times W}$, Rendered images $I_{r}$, Style images $I_{s}$
0: Scene-level style loss $\mathcal{L}_{\text{scene}}$
1:for $i=1$ to $B$ do
2:  Extract style features $F_{s}\leftarrow\text{VGG}(I_{s}[i])$
3:  Extract batched content features $F_{c}\leftarrow\text{VGG}(I_{c}[i,:])$
4:  Extract batched rendered features $F_{r}\leftarrow\text{VGG}(I_{r}[i,:])$
5:  Content loss:
$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
6:  Style loss initialization: $\mathcal{L}_{s}\leftarrow 0$
7:  for $k=1$ to 55 do
8:   Reshape $F_{r}^{(k)}$ into $(B,S,C,H,W)$
9:   $(\mu_{r},\sigma_{r})\leftarrow\text{MeanStd}(F_{r}^{(k)})$
10:   $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
11:   $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{r}-\mu_{s}\|_{2}^{2}+\|\sigma_{r}-\sigma_{s}\|_{2}^{2}$
12:  end for
13:end for
14:return $\mathcal{L}_{\text{scene}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
Algorithm 3 3D Voxel-based AdaIN Style Loss
0: Rendered features $F_{r}$, Style features $F_{s}$, Content features $F_{c}$, 3D points $P\in\mathbb{R}^{B\times S\times H\times W\times 3}$, Confidence $C$, Valid mask $M$
0: 3D style loss $\mathcal{L}_{\text{3d}}$
1:for $i=1$ to $B$ do
2:  for $k=1$ to 55 do
3:   Resize $(C,M,P)$ to match the resolution of $F_{r}^{(k)}$
4:   for $s=1$ to $S$ do
5:    $\mathcal{V}\leftarrow\text{Voxelize}(F_{r}^{(k)}[s],P[s],C[s],M[s])$
6:    $(\mu_{v},\sigma_{v})\leftarrow\text{MeanStd}(\mathcal{V})$
7:    $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
8:    $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{v}-\mu_{s}\|_{2}^{2}+\|\sigma_{v}-\sigma_{s}\|_{2}^{2}$
9:   end for
10:  end for
11:  Content loss:
$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
12:end for
13:return $\mathcal{L}_{\text{3d}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
A.2 Efficiency Scaling
(a) Time per batch vs. # views per batch
(b) GPU memory usage vs. # views per batch
Figure 6: Scaling of inference time and GPU memory with the number of views per batch. Time is averaged over 100 iterations, and memory is reported as the peak allocation, without rendering cost.
We further analyze the efficiency of Stylos by varying the number of views processed per batch. The results in Fig. 6 indicate a clear scaling pattern. For very small batches, such as 1–2 views, the average runtime per batch is low, around 0.3 s, but the measurements fluctuate heavily due to overheads and GPU scheduling. This makes single-view or extremely small-batch inference unstable, even though the raw latency is minimal.
As the batch size increases, runtime grows roughly linearly, reaching 0.75 s at 16 views and 4.83 s at 64 views. Importantly, once the batch size exceeds 4–8 views, the runtime becomes much more stable, showing only minor variation across repeated runs.
Memory usage shows a similarly consistent trend: GPU memory rises steadily from 4.16 GB at 1 view to 10.42 GB at 64 views. The growth is close to linear with respect to batch size, which makes resource requirements easy to estimate when scaling to larger workloads.
A.3 Additional Visual Results
Figure 7-16 presents additional qualitative results of Stylos on diverse style–content pairs. Across different styles and object categories, Stylos produces visually pleasing stylizations that respect the input content geometry while transferring the target artistic appearance. These examples further demonstrate the versatility and robustness of our approach beyond the main results shown in the paper.
Figure 7: Additional visual results of Stylos on diverse style–content pairs on Tanks and Temples Dataset (Knapitsch et al. (2017a)). Stylos consistently produces visually pleasing stylizations that preserve content geometry while transferring the target style.
Figure 8: Additional visual results of Stylos on the CO3D dataset. The top left is the style image while the rest of the top row are the rendered images. The mid row are the rendered images conditioned on the mid left imamge. The bottom row is the content inputs.
Figure 9: Additional visual results of Stylos on diverse style–content pairs.
Figure 10: Additional visual results of Stylos on diverse style–content pairs.
Figure 11: Additional visual results of Stylos on diverse style–content pairs.
Figure 12: Additional visual results of Stylos on diverse style–content pairs.
Figure 13: Additional visual results of Stylos on diverse style–content pairs.
Figure 14: Additional visual results of Stylos on the DL3DV-10K dataset. The top left is the style image while the rest of the top row are the rendered images. The mid row are the rendered images conditioned on the mid left imamge. The bottom row is the content inputs.
Figure 15: Additional visual results of Stylos on diverse style–content pairs.
Figure 16: Additional visual results of Stylos on diverse style–content pairs.

### v2

@@
 With the addition of a voxel-based 3D style loss that aligns aggregated scene features to style statistics, Stylos enforces view-consistent stylization while preserving geometry.
-Experiments across multiple datasets demonstrate that Stylos delivers high-quality zero-shot stylization, highlighting the effectiveness of global style–content coupling, the proposed 3D style loss, and the scalability of our framework from single view to large-scale multi-view settings.
-Our codes will be available at https://github.com/HanzhouLiu/StylOS.
+Experiments across multiple datasets demonstrate that Stylos delivers high-quality zero-shot stylization, highlighting the effectiveness of global style-content coupling, the proposed 3D style loss, and the scalability of our framework from single view to large-scale multi-view settings.
+Our codes are available at https://github.com/HanzhouLiu/Stylos.
+footnotetext: * Equal Contribution
 1 Introduction
@@
 Stylos employs a shared Transformer backbone with two pathways: content and style images are projected into a shared feature space, where content retains self-attention for geometric reasoning, and style is injected via global cross-attention to condition aesthetics.
-Notably, the global cross-attention also benefits the preservation of geometric details.
+Notably, the global cross-attention also helps preserve fine visual details in the Gaussian splatting renderings.
 Geometry-related attributes, such as depth, camera intrinsics, and extrinsics, are derived from backbone features, whereas style conditioning guides the prediction of color coefficients.
@@
 Implicit neural representations like NeRF (Mildenhall et al., 2020) improve geometric consistency, inspiring a range of stylization methods based on stylized view supervision or auxiliary networks (Nguyen-Phuoc et al., 2022; Zhang et al., 2022; Chiang et al., 2022; Huang et al., 2022).
-StyleRF (Liu et al., 2023) further enhances fidelity and enables zero-shot transfer by applying style transformations in radiance-field feature space.
-More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as an efficient alternative to NeRF.
-StyleGaussian (Liu et al., 2024) extends 3DGS with efficient feature rendering to enable real-time stylization, while subsequent works refine 3DGS for specialized settings and efficiency improvements (Mei et al., 2024; Zhang et al., 2024; Yu et al., 2024; Galerne et al., 2025).
-In parallel, Styl3R (Wang et al., 2025b) disentangles geometry and style using separate encoders, but is restricted to 2–8 input views.
-In contrast, Stylos follows the shared-backbone paradigm of AdaIN (Huang & Belongie, 2017), leveraging VGGT for unified feature extraction, and scales from a single to hundreds of input views.
+StylizedNeRF (Huang et al., 2022) addresses the domain gap between style images and NeRF by introducing a mutual learning framework.
+StyleRF (Liu et al., 2023) further enables zero-shot transfer by applying style transformations in radiance-field feature space.
+However, training NeRF for scene reconstruction is typically computationally expensive, making NeRF-based 3D stylization far from real-time.
+More recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as an efficient alternative to NeRF for novel view synthesis.
+Subsequent works further refine 3DGS for efficient stylization (Mei et al., 2024; Zhang et al., 2025; Kovács et al., 2024; Yu et al., 2024; Galerne et al., 2025).
+StyleGaussian (Liu et al., 2024) extends 3DGS with efficient feature rendering to enable real-time zero-shot stylization.
+However, despite being significantly faster to optimize than NeRF, 3DGS-based approaches still require per-scene fitting, which poses a fundamental barrier to achieving truly real-time 3D stylization.
+Styl3R (Wang et al., 2025b), the closest contemporaneous related work to ours, introduces a feedforward framework for instant 3D stylized reconstruction.
+While effective and efficient, Styl3R is primarily designed for 2–8 input views and does not specifically target strong multi-view consistency among the rendered stylization results.
 Underlying these architectures, style transfer objectives play a crucial role.
@@
 To this end, we extend statistics-based style losses into voxel space, introducing a 3D-aware objective that aligns feature distributions after multi-view fusion.
-This formulation complements perceptual and CLIP-based losses, enforcing geometry-aware and view-consistent stylization.
 Figure 1: Architecture overview.
@@
 where the scene representation $G=\{(p_{m},\,c_{m})\}_{m=1}^{M}$ is parameterized by $M$ anisotropic 3D Gaussians.
-Each Gaussian has geometry attributes $p_{m}=(\mu_{m},\,\alpha,\,r_{m},\,s_{m})$, comprising 3D position $\mu_{m}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}_{+}$, orientation quaternion $r_{m}\in\mathbb{R}^{4}$, and anisotropic scale $s_{m}\in\mathbb{R}^{3}$, as well as a style-dependent color embedding $c_{m}\in\mathbb{R}^{3\times(k+1)^{2}}$ represented with spherical harmonics of degree $k$.
+Each Gaussian has geometry attributes $p_{m}=(\mu_{m},\,\alpha_{m},\,r_{m},\,s_{m})$, comprising 3D position $\mu_{m}\in\mathbb{R}^{3}$, opacity $\alpha\in\mathbb{R}^{+}$, orientation quaternion $r_{m}\in\mathbb{R}^{4}$, and anisotropic scale $s_{m}\in\mathbb{R}^{3}$, as well as a style-dependent color embedding $c_{m}\in\mathbb{R}^{3\times(k+1)^{2}}$ represented with spherical harmonics of degree $k$.
 In addition, Stylos jointly predicts camera parameters $\{g_{i}\in\mathbb{R}^{9}\}_{i=1}^{N}$ for each input view, leveraging pose cues estimated by the VGGT backbone.
 3.2 Network Architecture
-Architecture Overview.
-At its core, Stylos employs a geometry backbone that alternates frame-wise and global attention to process multiple input views and infer geometry-related parameters such as positions, scales, orientations, and opacities.
+Stylos employs a geometry backbone that alternates frame-wise and global attention to process multiple input views and infer geometry-related parameters such as positions, scales, orientations, and opacities.
 To enable stylization, we introduce a dedicated conditioning branch, the Style Aggregator, which fuses content and style features through cross-attention and predicts style-aware color embeddings for each Gaussian.
@@
 3.2.2 Style Aggregator
-The Style Aggregator extends the geometric backbone by adapting its attention layers from self-attention to cross-attention.
-Both content views and the style image are embedded with a shared patch encoder, ensuring that their tokens lie in a common feature space.
-In this design, content tokens serve as queries while style tokens provide keys and values, so that style information conditions the token representations later used for predicting Gaussian colors.
-Geometry-related parameters, in contrast, are still predicted from the backbone pathway and remain unaffected by style conditioning.
-Formally, let $T\in\mathbb{R}^{B\times N\times K\times C}$ denote the content tokens from $N$ views, each with $K$ patches and $C$ channels, and let $S\in\mathbb{R}^{B\times K\times C}$ represent the style tokens.
-Let $T_{i}$ be the token set of the $i$-th view, and let $\mathrm{Attn}(Q,KV)$ denote a standard cross-attention operation.
-Depending on how queries are formed, we distinguish three style–content coupling mechanisms,
-(Frame-only): $T^{\text{frame}}_{i}=\mathrm{Attn}(T_{i},S),$ $Q=T_{i},\;\;K=V=S,$ (1)
-(Global-only): $T^{\text{global}}=\mathrm{Attn}(\{T_{i}\}_{i=1}^{N},S),$ $Q=\{T_{i}\}_{i=1}^{N},\;\;K=V=S,$ (2)
-(Hybrid): $T^{\text{hybrid}}=[\,\{T^{\text{frame}}_{i}\}_{i=1}^{N}\,\|\,T^{\text{global}}\,],$ concatenation along channel dim. (3)
-Frame-only injects style independently per view, global-only enforces a shared conditioning across all views, and hybrid concatenates the frame-level and global-level outputs along the channel dimension, yielding $2\times C$-dimensional features.
-In our design, Stylos adopts the global-only configuration, implemented with CrossBlock layers where $\{T_{i}\}_{i=1}^{N}$ serve as queries and $S$ provides keys and values.
-The resulting features are passed to the prediction heads described in Sec.
-3.2.3.
+The Style Aggregator builds on the geometric backbone inherited from VGGT, which is composed of standard Transformer $\mathrm{Block}$ containing self-attention followed by feed-forward layers.
+To enable style conditioning, we replace this standard block with an adapted $\mathrm{CrossBlock}$, which inserts a cross-attention operation between the self-attention and MLP stages (Deng et al., 2022).
+In $\mathrm{CrossBlock}$, content tokens serve as queries ($\mathcal{Q}$) while style tokens provide keys and values ($\mathcal{KV}$).
+This structure allows the content representations to be refined by their own spatial context (via the preserved self-attention) while being explicitly conditioned on the target style (via the inserted cross-attention) before the final feed-forward projection.
+Content–Style Coupling.
+Let $\mathcal{Q}_{b,v}\in\mathbb{R}^{L_{q}\times C}$ denote the content tokens extracted from the $v$-th view of the $b$-th sample, and let $\mathcal{KV}_{b}\in\mathbb{R}^{L_{kv}\times C}$ denote the style tokens.
+We use $\mathrm{CrossBlock}(\mathcal{Q},\mathcal{KV})$ to denote our block operator that updates $\mathcal{Q}$ using $\mathcal{KV}$ as key–value pairs.
+Depending on how the query set is formed, we consider two primary topological strategies.
+(1) Frame cross attention. Each view independently interacts with the style tokens,
+$\widetilde{\mathcal{Q}}_{b,v}=\mathrm{CrossBlock}(\mathcal{Q}_{b,v},\,\mathcal{KV}_{b}).$ (1)
+Since the internal self-attention of the block operates only on $\mathcal{Q}_{b,v}$, this strategy preserves view-specific geometric structure but prevents information propagation between different views.
+(2) Global cross attention.
+We first concatenate all views to obtain a global sequence $\mathcal{Q}_{b}^{\mathrm{global}}\in\mathbb{R}^{VL_{q}\times C}$ by $\mathcal{Q}_{b}^{\mathrm{global}}=\mathrm{Concat}_{v=1}^{V}\mathcal{Q}_{b,v}$, and perform the block operation simultaneously,
+$\widetilde{\mathcal{Q}}^{\mathrm{global}}_{b}=\mathrm{CrossBlock}(\mathcal{Q}^{\mathrm{global}}_{b},\,\mathcal{KV}_{b}).$ (2)
+We then reshape the updated tokens back to per-view tensors, $\widetilde{\mathcal{Q}}_{b,v}$, by splitting $\widetilde{\mathcal{Q}}^{\mathrm{global}}_{b}$ along the sequence dimension.
+This approach enables long-range reasoning: the internal self-attention ensures multi-view geometric consistency, while the cross-attention broadcasts style information globally.
+(3) Hybrid cross attention.
+We first apply frame attention to refine each view independently as in Eq.
+1, and then perform global attention by concatenating the resulting $\widehat{\mathcal{Q}}_{b,v}$ across views and applying the same operator as in Eq.
+2.
+This hierarchical coupling combines per-view refinement with multi-view aggregation.
 3.2.3 Prediction Heads
@@
 Auxiliary Heads, Adapter, and Voxelization.
-We employ the existing VGGT camera head to estimate camera intrinsics and extrinsics, and a depth head to predict scene geometry cues, which are unprojected into 3D anchors for Gaussian placement (Kerbl et al., 2023; Ren et al., 2024).
+We employ the existing VGGT camera head to estimate camera intrinsics and extrinsics, and a depth head to predict scene geometry cues, which are unprojected into 3D anchors for Gaussian placement (Kerbl et al., 2023; Ren et al., 2025).
 A Gaussian adapter then consolidates geometry and style outputs into a unified set of primitives $\{(p_{m},c_{m})\}_{m=1}^{M}$ for differentiable rendering.
@@
 $\mathcal{L}_{\text{stage2}}=\mathcal{L}_{\text{rec}}+\lambda_{\text{style}}\,\mathcal{L}_{\text{style}}^{3D}+\lambda_{\text{cnt}}\,\mathcal{L}_{\text{content}}+\lambda_{\text{clip}}\,\mathcal{L}_{\text{clip}}+\lambda_{\text{tv}}\,\mathcal{L}_{\text{TV}}.$
+In our experiments, we set $\lambda_{\text{distill}}=1.0$, and use $\{{\lambda_{\text{style}},\lambda_{\text{cnt}},\lambda_{\text{clip}},\lambda_{\text{tv}}}\}=\{{1.0,0.1,1.0,10.0}\}$
 3.4 Style Losses
-We denote by $F^{l}_{b,s}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the VGG feature map of the rendered (stylized) image for the $b$-th scene and $s$-th view at layer $l$, and by $S^{l}_{b}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the feature map of the style reference image.
-Building on the BN-statistics formulation (Li et al., 2017), which interprets channel-wise mean and variance as style descriptors, we progressively extend statistics matching beyond the standard 2D setting.
-This progression starts from independent image-level matching, moves to multi-view feature aggregation for enforcing cross-view consistency, and culminates in a voxel-space loss that directly constrains the fused 3D representation.
-Image-Level Style Loss. The simplest baseline aligns each rendered frame independently with the style reference,
-$\mathcal{L}^{\mathrm{img}}_{\mathrm{sty}}=\frac{1}{BS}\sum_{b=1}^{B}\sum_{s=1}^{S}\sum_{l}\alpha_{l}\Bigl(\|\mu(F^{l}_{b,s})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(F^{l}_{b,s})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (4)
-This encourages per-frame stylization but does not enforce multi-view consistency.
+We denote by $\mathcal{R}^{l}_{b,v}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the VGG feature map of the rendered image for the $b$-th scene and $v$-th view at layer $l$, and by $\mathcal{S}^{l}_{b}\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ the feature map of the single style image.
+Building on the Batch Normalization statistics method (Li et al., 2017), which interprets BN mean and variance as style descriptors, we progressively extend it matching beyond standard 2D settings.
+This progression starts from independent image-level matching, moves to multi-view feature aggregation that promotes global style coherence across views, and culminates in a voxel-space loss that directly constrains the fused 3D representation.
+Next, we formulate the three losses studied in this paper.
+Image-Level Style Loss. The simplest baseline aligns each rendered frame independently with the style reference, which encourages per-frame stylization but does not enforce multi-view consistency,
+$\mathcal{L}^{\mathrm{img}}_{\mathrm{sty}}=\frac{1}{BV}\sum_{b=1}^{B}\sum_{v=1}^{V}\sum_{l=1}^{5}\alpha_{l}\Bigl(\|\mu(\mathcal{R}^{l}_{b,v})-\mu(\mathcal{S}^{l}_{b})\|_{2}^{2}+\|\sigma(\mathcal{R}^{l}_{b,s})-\sigma(\mathcal{S}^{l}_{b})\|_{2}^{2}\Bigr),$ (3)
+where $\alpha_{l}$ denotes the weight at each level $l$, $\mu$ is the mean and $\sigma$ is the standard deviation.
 Scene-Level Style Loss.
-To encourage consistent stylization across multiple views, we concatenate the per-view features $\{F^{l}_{b,s}\}_{s=1}^{S}$ along the spatial dimension to form $\widetilde{F}^{l}_{b}$, and compute statistics on this aggregated map,
-$\mathcal{L}^{\mathrm{scn}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l}\alpha_{l}\Bigl(\|\mu(\widetilde{F}^{l}_{b})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(\widetilde{F}^{l}_{b})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (5)
-This step enforces cross-view consistency but still operates in 2D feature space.
-3D Style Loss.
-Finally, we fuse multi-view features into a voxel grid using differentiable unprojection, where features from different views are accumulated into spatial bins of a discretized 3D volume.
-Let $G^{l}_{b}$ denote the voxelized features for the $b$-th scene at layer $l$.
+To introduce consistency into multi-view 3D stylization, we concatenate the per-view features $\{\mathcal{R}^{l}_{b,v}\}_{v=1}^{V}$ along the spatial dimension to form $\widetilde{\mathcal{R}}^{l}_{b}$, and calculate statistics on this aggregated map.
+However, this method still operates in 2D feature space and is computed by,
+$\mathcal{L}^{\mathrm{scn}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l=1}^{5}\alpha_{l}\Bigl(\|\mu(\widetilde{\mathcal{R}}^{l}_{b})-\mu(\mathcal{S}^{l}_{b})\|_{2}^{2}+\|\sigma(\widetilde{\mathcal{R}}^{l}_{b})-\sigma(\mathcal{S}^{l}_{b})\|_{2}^{2}\Bigr).$ (4)
+Algorithm 1 3D Voxel-based AdaIN Style Loss
+0: Rendered features $\mathcal{R}$, Style features $\mathcal{S}$, 3D points $\mathcal{P}$, Confidence $\mathcal{C}$, Valid mask $\mathcal{M}$
+0: 3D style loss $\mathcal{L}_{sty}^{\text{3D}}$
+1: Initialize $\mathcal{L}_{sty}^{3D}\leftarrow 0$
+2:for $b=1$ to $B$ do
+3:  for $l=1$ to 55 do
+4:   Resize $(\mathcal{C,M,P})$ to match the resolution of $\mathcal{R}^{l}_{b}$
+5:   // Next, fuse all views into a single voxel grid
+6:   $\mathcal{G}_{b}^{l}\leftarrow\text{VoxelizeAndFuse}\left(\{\mathcal{R}^{l}_{b,v}\}_{v=1}^{V},\{\mathcal{P}_{b,v}\}_{v=1}^{V},\{\mathcal{C}_{b,v}\}_{v=1}^{V},\{\mathcal{M}_{b,v}\}_{v=1}^{V}\right)$
+7:   // Next, compute the BN statistics (mean and standard deviation) of two feature maps
+8:   $(\mu_{g},\sigma_{g})\leftarrow\text{MeanStd}(\mathcal{G}_{b}^{l})$, $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(\mathcal{S}_{b}^{l})$
+9:   $\mathcal{L}_{sty}^{3D}\leftarrow\mathcal{L}_{sty}^{3D}+\alpha_{l}\|\mu_{g}-\mu_{s}\|_{2}^{2}+\alpha_{l}\|\sigma_{g}-\sigma_{s}\|_{2}^{2}$
+10:  end for
+11:end for
+12:return $\mathcal{L}_{sty}^{3D}=\lambda_{sty}\mathcal{L}_{sty}^{3D}$
+Voxel-level 3D Style Loss.
+Finally, to further enforce multi-view consistency, we fuse multi-view features into a voxel grid using differentiable unprojection, where features from different views are accumulated into spatial bins of a discretized 3D volume.
+Let $\mathcal{G}^{l}_{b}$ denote the voxelized features for the $b$-th scene at layer $l$.
 We then compute style statistics directly in voxel space,
-$\mathcal{L}^{\mathrm{3D}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l}\alpha_{l}\Bigl(\|\mu(G^{l}_{b})-\mu(S^{l}_{b})\|_{2}^{2}+\|\sigma(G^{l}_{b})-\sigma(S^{l}_{b})\|_{2}^{2}\Bigr).$ (6)
-By operating on voxelized features, this loss explicitly encodes geometry and enforces style consistency across both views and the underlying 3D structure.
-Figure 2: CO3D pizza scene comparing different style–content cross-attention strategies.
+$\mathcal{L}^{\mathrm{3D}}_{\mathrm{sty}}=\frac{1}{B}\sum_{b=1}^{B}\sum_{l=1}^{5}\alpha_{l}\Bigl(\|\mu(\mathcal{G}^{l}_{b})-\mu(\mathcal{S}^{l}_{b})\|_{2}^{2}+\|\sigma(\mathcal{G}^{l}_{b})-\sigma(\mathcal{S}^{l}_{b})\|_{2}^{2}\Bigr).$ (5)
+By operating on voxelized features, this loss explicitly encodes geometry and enforces style consistency across both views and the underlying 3D structure. We provide pseudo codes in Algorithm 1.
+4 Experiment
+Datasets.
+We evaluate cross-category generalization on the CO3D (Reizenstein et al., 2021) dataset by training on 17 categories and testing on 3 held-out ones, and cross-scene generalization by training on the full DL3DV-10K (Ling et al., 2024) and testing on Tanks & Temples (Knapitsch et al., 2017).
+Style images are provided by WikiArt (WikiArt, 2025) and DELAUNAY (Gontier et al., 2023), with 50 diverse style images reserved as unseen styles that are never used during training.
+Baselines.
+We compare Stylos with recent 3D stylization models.
+The baselines include, (1) a per-scene training and zero-shot method StyleGaussian (Liu et al., 2024), (2) per-scene and per-style optimization approaches G-Style (Kovács et al., 2024), StylizedGS (Zhang et al., 2025), and SGSST (Galerne et al., 2025), and (3) the closest related work Styl3R (Wang et al., 2025b).
+Evaluation Metrics.
+Our evaluation covers three aspects.
+(1) To assess geometry reconstruction in the style-free setting, we report PSNR, SSIM, and LPIPS (Zhang et al., 2018) between content views and predictions, conditioned on the original first frame of each scene as the style input.
+(2) To measure stylization consistency, following prior work (Chiang et al., 2022), we compute LPIPS and RMSE in both short-range and long-range settings.
+(3) To further evaluate stylization quality, we report ArtScore (Chen et al., 2024), a recent metric specifically designed for reference-free evaluation of artness in generated images, and ArtFID (Wright & Ommer, 2022; Chung et al., 2024).
+4.1 Ablation Study
+In this section, we analyze the impact of each design choice on the CO3D dataset. To ensure fairness, all variants are trained with identical configurations in each experimental round.
 […]
@@
 21.68 0.7043 0.2684 20.57 0.6177 0.3110 22.09 0.7362 0.3125
-4 Experiment
-Datasets.
-We evaluate cross-category generalization on the CO3D (Reizenstein et al., 2021) dataset by training on 17 categories and testing on 3 held-out ones, and cross-scene generalization by training on the full DL3DV-10K (Ling et al., 2024) and testing on Tanks and Temples (Knapitsch et al., 2017b).
-Style supervision is provided by WikiArt (WikiArt, ), with 50 images reserved for unseen-style evaluation.
-Baselines.
-We compare with two representative open-source baselines, StyleRF (Liu et al., 2023) and StyleGS (Liu et al., 2024), using their released per-scene trained weights against our single DL3DV-trained model.
-Concurrent to Stylos, the other instant stylization solution Styl3R (Wang et al., 2025b) was proposed, but its official codes are not yet available at the time of our submission.
-We plan to provide comparisons once they are released.
-Evaluation Metrics.
-Our evaluation covers three aspects.
-(1) To assess geometry reconstruction in the style-free setting, we report PSNR, SSIM, and LPIPS (Zhang et al., 2018) between content views and predictions, conditioned on the original first frame of each scene as the style input.
-(2) To measure stylization consistency, following prior work (Chiang et al., 2022), we compute LPIPS and RMSE in both short-range (consecutive frames $t$ and t1t{-}1) and long-range (distant frames $t$ and t7t{-}7) settings.
-(3) To further evaluate stylization quality, we report ArtScore (Chen et al., 2024), a recent metric specifically designed for reference-free evaluation of artness in generated images.
-4.1 Ablation Study
-In this section, we analyze the impact of each design choice on the CO3D dataset. To ensure fairness, all variants are trained with identical configurations in each experimental round.
+Figure 2: CO3D pizza scene comparing different style–content cross-attention strategies.
 Style-Content Cross Attention.
-As reported in Table 1, restricting cross-attention to the frame level yields only moderate performance, with a PSNR of 20.9 dB on skateboard and 19.7 dB on pizza.
-The hybrid variant, which combines frame-level and global attention, provides observable improvement on donut with 21.4 dB, but brings limited gains on the other categories and overall underperforms the global-only setup.
-By contrast, applying cross-attention at the global level consistently achieves the best results, reaching a PSNR of 21.7 dB on skateboard, 20.6 dB on pizza, and 22.1 dB on donut.
+Take the pizza scene for example.
+Table 1 shows that the baseline method, combining frame-level and global attention, yields a PSNR of 19.78 dB and LPIPS of 0.3326.
+The frame variant produces slightly lower PSNR and higher LPIPS.
+By contrast, applying cross-attention at the global level significantly boosts the PSNR value by 0.79 dB and achieves consistently improved metrics.
 Qualitative comparisons in Fig.
-2 highlight these differences on the pizza example.
-Frame-only attention produces a poorly defined crust with noticeable artifacts, the hybrid variant recovers part of the structure but leaves the box edges blurred and unstable, while global attention delivers the most faithful result, with clear toppings and a well-preserved crust boundary.
+2 highlight these differences.
+Frame-only attention produces a poorly defined crust with noticeable artifacts, the hybrid variant recovers part of the structure but leaves the box edges blurred, while global attention delivers the most faithful result, with clear toppings and a well-preserved crust boundary.
 Figure 3: Comparison of style losses on unseen donut, skateboard, and pizza scenes from the CO3D dataset.
 Both scene and 3D style losses yield cleaner stylized textures compared to image-level matching, while the 3D loss further conveys a stronger sense of 3D geometry.
-Table 2: Consistency comparison of different style loss designs on CO3D. The 3D loss achieves the strongest short-range and long-range consistency.
-Style Loss Short-range Long-range
+We encourage readers to Appendix Fig.
+8-10 for more visual comparisons under varing scenes and styles.
+Table 2: Comparison of consistency and artistic quality among different style loss designs on CO3D. The frame stride is set to 3, and 15 held-out scenes are randomly selected for evaluation.
+[…]
 $downarrow LPIPS\downarrow RMSE\downarrow$
-Image loss (Eq. 4) 0.046 0.035 0.147 0.135
-Scene loss (Eq. 5) 0.048 0.037 0.147 0.133
-3D loss (Eq. 6) 0.046 0.033 0.140 0.128
+Image loss (Eq. 3) 0.048 0.038 0.157 0.142 4.78
+Scene loss (Eq. 4) 0.047 0.036 0.156 0.148 9.12
+3D loss (Eq. 5) 0.047 0.034 0.153 0.142 9.15
 Image vs.
@@
 3D Style Losses.
-As shown in Table 2, the image-level loss achieves short-range LPIPS of 0.046 and RMSE of 0.035, but long-range consistency rises to 0.147 and 0.135.
-The scene-level loss slightly improves long-range RMSE to 0.133 but degrades short-range performance to 0.048 and 0.037, indicating that simple 2D aggregation reduces some long-range inconsistencies but compromises short-range accuracy.
-In contrast, the 3D loss applies statistics after voxel-space fusion, matching the best short-range LPIPS of 0.046, further lowering short-range RMSE to 0.033, and improving long-range scores to 0.140 and 0.128.
-These results indicate that enforcing style consistency in a geometry-aligned space yields sharper boundaries and more coherent textures across views.
-As further illustrated in Fig.
-3, image-level supervision often fails to faithfully transfer style.
-For example, the donut lacks stylization and the wooden surface is not well synthesized, whereas scene- and 3D-level losses successfully apply the target style.
-Among them, the 3D style loss produces sharper boundaries and a stronger sense of 3D geometry, as observed in the skateboard and pizza examples, leading to the most coherent multi-view stylization.
+As shown in Table 2, both scene-level and 3D style losses clearly outperform the image-level baseline in terms of artistic quality.
+Overall, the 3D losses exhibit superior and more stable consistency performance.
+These results indicate that enforcing style consistency in a voxel-level yields more coherent stylized textures across views.
+We observe that the image-level supervision may fail to transfer style.
+For example, in Fig.
+3, the donut surface is not well synthesized with the expected style, whereas the scene and 3D losses successfully apply the target style.
+The 3D style loss produces sharper boundaries and a stronger sense of 3D geometry, as observed in the skateboard and pizza examples, leading to the most coherent multi-view stylization.
+We release these model weights to support reproducibility and further analysis.
 From a Single View to Dozens of Views.
-While Stylos can process hundreds of content inputs, we qualitatively observe a gradual decrease in visual quality once the number of views per batch (denoted as “# views / batch”) exceeds 32.
+While Stylos can process up to dozens (even hundreds) of views, we observe a gradual decrease in visual quality once the number of views per batch (denoted as “# views / batch”) exceeds 32.
 As shown in Fig.
-4, very small batches (1 view) fail to synthesize parts of the scene such as the bench, while overly large batches (64 views) introduce edge artifacts on the tall building, potentially due to unstable Gaussian representations and the gap from training settings (no more than 24 views).
+4, very small batches (1 view) fail to synthesize parts of the scene like the bench, while overly large batches (64 views) introduce edge artifacts on the tall building, potentially due to the gap from our training settings (no more than 24 views).
+Additionally, we provide A.2 Fig.
+7 to show the efficiency trend with respect to the number of views.
 Figure 4: Effect of varying # views / batch on the Lighthouse scene from Tanks and Temples.
@@
 Short-range consistency
-StyleRF 2023 0.075 0.065 0.060 0.062 0.039 0.042
-StyleGS 2024 0.033 0.038 0.031 0.034 0.038 0.037 0.069 0.061
+StyleGaussian  2024 0.033 0.038 0.031 0.034 0.038 0.037 0.069 0.061
+G-Style  2024 0.042 0.052 0.032 0.035 0.038 0.034 0.066 0.070
+SGSST  2025 0.038 0.047 0.039 0.047 0.044 0.049 0.084 0.090
+Styl3R 2025b 0.061 0.036 0.066 0.040 0.105 0.067
 Stylos (ours) 0.030 0.026 0.028 0.021 0.035 0.024 0.047 0.044
 Long-range consistency
-StyleRF 2023 0.116 0.110 0.176 0.154 0.096 0.092
-StyleGS 2024 0.067 0.072 0.086 0.077 0.091 0.091 0.177 0.141
+StyleGaussian  2024 0.067 0.072 0.086 0.077 0.091 0.091 0.177 0.141
+G-Style  2024 0.098 0.120 0.095 0.093 0.104 0.095 0.180 0.175
+SGSST  2025 0.087 0.108 0.119 0.120 0.130 0.128 0.221 0.222
+Styl3R  2025b 0.116 0.100 0.147 0.143 0.146 0.145
 Stylos (ours) 0.051 0.056 0.074 0.069 0.083 0.082 0.139 0.134
+  • StylizedGS (Zhang et al., 2025) is not included in quantitative comparisons due to its multiple failure cases observed on our test styles.
+Nevertheless, its quantitative results are reported in A.4 Table 5 and Table 6 for readers’ reference.
+Reproduced results are available at https://github.com/HanzhouLiu/Stylos.
 4.2 Comparison with State-of-the-Art
-For StyleRF (Liu et al., 2023) and StyleGS (Liu et al., 2024), we directly evaluate with their released per-scene weights.
-In contrast, our Stylos is trained once on the full DL3DV dataset (Ling et al., 2024) and tested in a zero-shot manner, without ever seeing the images of these scenes.
+To retrain StyleGaussian (Liu et al., 2024), G-Style (Kovács et al., 2024), StylizedGS (Zhang et al., 2025), and SGSST (Galerne et al., 2025), we strictly follow their released codes.
+Since the pretrained-weights of StyleGaussian (Liu et al., 2024) and Styl3R (Wang et al., 2025b) are publicly available, we use them to generate the visual results.
+Stylos is trained once on DL3DV (Ling et al., 2024), and tested in a zero-shot manner without prior knowledge of either the scenes or styles.
 Quantitative Evaluation.
-According to Table 3, Stylos achieves consistently lower LPIPS and RMSE than StyleGS across both short- and long-range view pairs.
-This demonstrates that Stylos provides stronger cross-view consistency.
-Furthermore, Table 4 reports that Stylos achieves the highest ArtScore across nearly all four Tanks and Temples scenes (Knapitsch et al.
-(2017a)), outperforming both StyleRF and StyleGS.
+As shown in Table 3, Stylos achieves strong and stable consistency scores, ranking the first across all consistency metrics and all four scenes.
+This indicates that Stylos provides markedly improved cross-view stylization consistency.
+Furthermore, Table 4 shows that Stylos attains either the best or second-best artistic metric values, as rated by ArtScore and ArtFID, on all four Tanks and Temples scenes (Knapitsch et al., 2017), while maintaining the fastest stylization speed.
+Overall, Stylos demonstrates a favorable balance between visual quality, consistency, and efficiency, suggesting its potential for practical real-time 3D stylization.
+Table 4: Stylization quality, as measured by ArtScore and ArtFID (abbreviated as Score and FID respectively), and stylization time comparisons with recent 3D stylization models.
+Stylos achieves consistently favorable metric scores across the four scenes.
+Additionally, we follow StyleID (Chung et al., 2024) and calculate additional metrics as reported in A.4 Table 5 and Table 6.
 […]
-Method Train Truck M60 Garden
-StyleRF 2023 7.71 4.12 8.48
-StyleGS 2024 0.78 5.76 8.63 9.38
-Stylos (ours) 9.50 9.70 9.37 9.34
-Figure 5: Visual comparison between Stylos and recent per-scene 3D stylization baselines.
+[…]
+StyleGaussian 2024 0.78 52.79 5.76 44.93 8.63 47.48 9.38 41.14 165 m*
+G-Style 2024 9.52 23.24 9.67 22.15 9.73 22.36 8.98 25.76 14.7 m*
+SGSST 2025 1.84 38.24 5.34 32.34 5.26 38.73 4.89 33.54 35.2 m*
+Styl3R 2025b 2.94 34.11 2.96 29.86 4.09 38.28 0.16 s
+Stylos (ours) 9.50 26.40 9.70 28.71 9.37 27.44 9.34 28.06 0.05 s
+    For methods with per-scene fitting required, the training phase accounts towards stylization time (see A.4).
+  • As to single-forward 3D stylization approaches, training is not considered for stylization time.
+Figure 5: Visual comparisons between Stylos and recent 3D stylization approaches. Stylos successfully transfers diverse artistic styles to the scenes while preserving fine structural details.
 Qualitative Evaluation.
-Fig.
-5 shows that Stylos successfully transfers the target style while preserving scene geometry and delivering consistent textures.
-In contrast, StyleGS introduces noticeable geometric distortions and texture discontinues, while StyleRF tends to over-impose style patterns, negatively impacting the object structure.
-Overall, Stylos achieves the best trade-off between structural fidelity and faithful stylization.
+Taking the truck scene with the desert-town style (yellow foreground objects with blue backgrounds) in Fig.
+5 as an example, we observe that nearlly all existing approaches could reasonably produce structured results.
+However, their differences are discussed as follows.
+The zero-shot method StyleGaussian fail to generate color-consistent scenes, several portions of the truck retain undesired color blocks.
+Per-scene and per-style fitting approaches such as G-Style, StylizedGS, and SGSST often fail to achieve complete style transfer, the truck door remaining predominantly blue instead of adopting the target yellow hue.
+The other feedforward solution Styl3R similarly struggles to propagate the correct yellow color throughout the truck.
+In contrast, Stylos successfully renders the truck in coherent yellow while preserving clean geometry, and correctly assigns blue tones to the distant background regions, producing an appearance that closely resembles the intended desert-town aesthetic.
+These observations underscore the strength of Stylos in achieving both faithful style expression and robust 3D structural preservation.
+4.3 Evaluation of Multi-Style Blending and Stylization Control
+We validate the model’s controllable stylization capability through two interpolation experiments.
+First, by linearly interpolating between the embeddings of two distinct style images, we observe smooth, coherent transitions that demonstrate the model’s natural support for multi-style blending (see Fig.
+6).
+Second, because the model is trained to reconstruct the original appearance when the style image matches the content image, we interpolate between the reconstruction (content) embedding and a stylized embedding Fig.
+6.
+This produces a continuous spectrum of outputs with gradually increasing stylization strength, confirming that our approach enables fine-grained, post-inference control over the content–style trade-off.
+Together, these results show that our method supports both multi-style fusion and adjustable stylization without any additional optimization.
+Figure 6: Evaluation of multi-style blending and controllable stylization. Top: interpolation between two style embeddings. Bottom: interpolation between the content embedding and a style embedding.
 5 Conclusion
@@
 In the future, we aim to scale Stylos to support higher-resolution inputs while improving efficiency, paving the way for practical 3D content creation.
+6 Acknowledgement
+This research used both the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois, and the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois.
+Delta and DeltaAI are joint efforts of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
+The allocations on Delta and DeltaAI [allocation number CIS250529] were made by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
 7 Reproducibility Statement
@@
 3.
-We release detailed pseudo-code for all loss functions in Appendix A, and will release the full implementation upon acceptance.
-However, to support reproducibility, we provide detailed pseudo codes of the proposed style loss functions (image-level, scene-level, and 3D-level) in Appendix A.1.
+We provide detailed pseudo-code for the proposed 3D style loss function in Algo.
+1, and will release the full implementation upon acceptance.
 All datasets used are publicly available, as described in Sec.
@@
 These resources should allow researchers to reproduce and extend our findings.
-In addition, we provide extensive qualitative results in Appendix A.3, and analyze efficiency trends with respect to the number of views per batch in Appendix A.2.
-Efficiency experiments are averaged over 100 iterations per setting.
+In Appendix A.4, we clarify the training and inference details of the state-of-the-art methods for reproducible comparisons.
+In addition, we provide extensive qualitative results in Appendix A.6, and analyze efficiency trends with respect to the number of views per batch in Appendix A.2, where efficiency experiments are averaged over 100 iterations per setting.
 Appendix A Appendix
-A.1 Style Loss Pseudo Codes
-We provide pseudo codes to the three variants for style losses introduced in the main paper, image-level, scene-level, and 3d style losses.
-This is for reproducibility purpose, and please follow the implementation in our codes which will be released in the near future.
-Algorithm 1 Image-level AdaIN Style Loss
-0: Content images $I_{c}\in\mathbb{R}^{B\times S\times C\times H\times W}$, Rendered images $I_{r}$, Style images $I_{s}$
-0: Image-level style loss $\mathcal{L}_{\text{img}}$
-1:for $i=1$ to $B$ do
-2:  Extract style features $F_{s}\leftarrow\text{VGG}(I_{s}[i])$
-3:  for $j=1$ to $S$ do
-4:   Extract content features $F_{c}\leftarrow\text{VGG}(I_{c}[i,j])$
-5:   Extract rendered features $F_{r}\leftarrow\text{VGG}(I_{r}[i,j])$
-6:   Content loss:
-$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
-7:   Style loss initialization: $\mathcal{L}_{s}\leftarrow 0$
-8:   for $k=1$ to 55 do
-9:    $(\mu_{r},\sigma_{r})\leftarrow\text{MeanStd}(F_{r}^{(k)})$
-10:    $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
-11:    $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{r}-\mu_{s}\|_{2}^{2}+\|\sigma_{r}-\sigma_{s}\|_{2}^{2}$
-12:   end for
-13:  end for
-14:end for
-15:return $\mathcal{L}_{\text{img}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
-Algorithm 2 Scene-level AdaIN Style Loss
-0: Content images $I_{c}\in\mathbb{R}^{B\times S\times C\times H\times W}$, Rendered images $I_{r}$, Style images $I_{s}$
-0: Scene-level style loss $\mathcal{L}_{\text{scene}}$
-1:for $i=1$ to $B$ do
-2:  Extract style features $F_{s}\leftarrow\text{VGG}(I_{s}[i])$
-3:  Extract batched content features $F_{c}\leftarrow\text{VGG}(I_{c}[i,:])$
-4:  Extract batched rendered features $F_{r}\leftarrow\text{VGG}(I_{r}[i,:])$
-5:  Content loss:
-$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
-6:  Style loss initialization: $\mathcal{L}_{s}\leftarrow 0$
-7:  for $k=1$ to 55 do
-8:   Reshape $F_{r}^{(k)}$ into $(B,S,C,H,W)$
-9:   $(\mu_{r},\sigma_{r})\leftarrow\text{MeanStd}(F_{r}^{(k)})$
-10:   $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
-11:   $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{r}-\mu_{s}\|_{2}^{2}+\|\sigma_{r}-\sigma_{s}\|_{2}^{2}$
-12:  end for
-13:end for
-14:return $\mathcal{L}_{\text{scene}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
-Algorithm 3 3D Voxel-based AdaIN Style Loss
-0: Rendered features $F_{r}$, Style features $F_{s}$, Content features $F_{c}$, 3D points $P\in\mathbb{R}^{B\times S\times H\times W\times 3}$, Confidence $C$, Valid mask $M$
-0: 3D style loss $\mathcal{L}_{\text{3d}}$
-1:for $i=1$ to $B$ do
-2:  for $k=1$ to 55 do
-3:   Resize $(C,M,P)$ to match the resolution of $F_{r}^{(k)}$
-4:   for $s=1$ to $S$ do
-5:    $\mathcal{V}\leftarrow\text{Voxelize}(F_{r}^{(k)}[s],P[s],C[s],M[s])$
-6:    $(\mu_{v},\sigma_{v})\leftarrow\text{MeanStd}(\mathcal{V})$
-7:    $(\mu_{s},\sigma_{s})\leftarrow\text{MeanStd}(F_{s}^{(k)})$
-8:    $\mathcal{L}_{s}\leftarrow\mathcal{L}_{s}+\|\mu_{v}-\mu_{s}\|_{2}^{2}+\|\sigma_{v}-\sigma_{s}\|_{2}^{2}$
-9:   end for
-10:  end for
-11:  Content loss:
-$\mathcal{L}_{c}\leftarrow\|F_{r}^{(4)}-F_{c}^{(4)}\|_{2}^{2}+\|F_{r}^{(5)}-F_{c}^{(5)}\|_{2}^{2}$
-12:end for
-13:return $\mathcal{L}_{\text{3d}}=\lambda_{c}\mathcal{L}_{c}+\lambda_{s}\mathcal{L}_{s}+\lambda_{\text{mse}}\|I_{r}-I_{c}\|^{2}$
+A.1 Implementation
+Implementation Details.
+For the geometry branch, we utilize VGGT as the backbone, comprising $L=24$ Alternating-Attention Transformer layers.
+Both the geometry transformer and the depth DPT head are initialized using pretrained VGGT weights, while all remaining layers are initialized randomly.
+The style aggregator similarly employs 24 Transformer layers, configured with content-global self-attention followed by content-style cross-attention.
+For voxelization, each output of Geo Head associates a confidence value, which is converted into a normalized intra-voxel weight via softmax over Gaussians within the same voxel; these weights determine each Gaussian’s contribution when aggregating voxel-level properties (opacity or color).
+The Gaussian Adapter maps the Geometry head and Style head predicted vector into explicit 3D Gaussian parameters.
+The module first splits the predicted vector into (i) isotropic scales, (ii) rotation quaternions, and (iii) degree-d spherical harmonic coefficients.
+Finally, we construct full covariance matrices using the predicted scale and rotation via the standard 3DGS covariance formulation.
+This yields the final set of Gaussian parameters: means, covariances, SH harmonics, opacities, scales, and normalized rotations.
+In the geometry training stage, the entire network is pre-trained.
+The geometry outputs, poses and depth, are guided by a frozen VGGT teacher.
+The DINO encoder is trained to capture color and style features.
+For stylization fine-tuning, we optimize only the style aggregator and the style head.
+We set the voxel size to 0.002 for differentiable voxelization.
+For both geometry pre-training and stylization fine-tuning, we train the model using the AdamW optimizer for 15k iterations.
+We employ a cosine learning-rate schedule with a peak learning rate of $2\times 10^{-4}$ and a 1k-iteration warmup.
+Data & Augmentation.
+Following Anysplat (Jiang et al., 2025), we randomly sample between 2 and 20 frames per clip, maintaining a fixed total of 20 frames per GPU.
+The input resolution is constrained to 448 pixels on the longest side, with the aspect ratio randomized between 0.5 and 1.0.
+Intrinsic augmentation is applied via random center-cropping (77%–100% of the original size) and random horizontal flipping.
+During Geometry Pretraining, we encourage the DINO encoder to better learn color and style tokens by occasionally replacing the content image with a style image and by applying color jitter to perturb its appearance, while keeping the original image as ground truth to supervise appearance-change learning.
+Loss Weights.
+Style weight $\lambda_{style}=1.0$, content-reconstruction weight $\lambda_{cnt}=0.1$, CLIP consistency weight $\lambda_{clip}=1.0$, and total-variance regularization weight $\lambda_{tv}=10.0$.
 A.2 Efficiency Scaling
@@
 The growth is close to linear with respect to batch size, which makes resource requirements easy to estimate when scaling to larger workloads.
-A.3 Additional Visual Results
-Figure 7-16 presents additional qualitative results of Stylos on diverse style–content pairs.
+A.3 Ablation Analysis
+Content-Style Coupling.
+Stylos consists of a style pathway which contributes to the color information for the predicted 3DGS representation, and a content pathway for the other 3DGS parameters.
+In the style pathway, we introduce a style aggregator with cross-attention to entangle the content and style tokens.
+Our training pipeline consists of two stages.
+The first stage focuses on 3D scene reconstruction, where a random content view is used as the style reference; Stage 2 performs actual stylization using a real artistic style image.
+When designing Stylos, it is therefore essential to verify that introducing content-style cross-attention in the style aggregator does not harm the reconstruction process.
+So, we compare three content–style coupling mechanisms (global, frame, and both) by calculating PSNR, SSIM and LPIPS between the rendered multi-view rgbs and their corresponding ground-thruth after Stage 1 training, using the first content view as the style reference.
+As shown in the Table 1, the global-only variant achieves the best reconstruction accuracy.
+This behavior is expected, because each view attends to the style tokens independently without interacting with the other views in frame attention.
+Treating a single view as the style reference for all views forces each view to match its patch-level representation to the same reference individually, which distorts the rendered images.
+Fig.
+2 shows that the frame attention significantly blurs the pizza box and toppings.
+Image-, Scene-, and Voxel-level Style Losses.
+In this section, we provide additional visual comparisons showing that the proposed scene and 3D style losses significantly outperform the baseline image style loss.
+Specifically, we observe that the scene- and 3D-level losses produce cleaner and more coherent textures, such as the wooden floor in Fig.
+8.
+As shown in Fig.
+9, the 3D style loss is also able to stylize a larger portion of the table surface compared with the image loss.
+In Fig.
+10, the seams on the tablecloth are well preserved under the 3D loss, while they become partially distorted when using the image or scene losses.
+We will release our implementations of all three style losses to facilitate further exploration by the community.
+Figure 8: Visual comparisons of different style losses on the CO3D skateboard scene.
+It is clear that the 3D style loss provides a stronger sense of 3D geometry and cleaner wooden floor textures evaluated on varying style images.
+Figure 9: Visual comparisons of different style losses on the CO3D pizza scene.
+It is clear that the 3D style loss provides stronger artistic stylization to the table surface, while the image style loss often fails to impose the expected styles to cover the whole table surface.
+Figure 10: Visual comparisons of different style losses on the CO3D donut scene.
+It is clear that the 3D style loss provides faithful stylization to the table cover while preserving the seams on the tablecloth compared to both the image style loss and the scene style loss.
+A.4 Reproducible Experiments
+Reproduce State-of-the-art Models.
+For StyleGaussian (Liu et al., 2024), SGSST (Galerne et al., 2025), StylizedGS (Zhang et al., 2025), and GStyle (Kovács et al., 2024), we use their officially released codebases and pretrained weights.
+We either run inference directly with the provided models or follow the authors’ instructions to train the models when necessary.
+All of them are evaluated on GH200 GPUs for fair comparisons.
+We notice that StylizedGS (Zhang et al., 2025) fail in serveral style cases, producing stylized images with pure colors with barely visible geomtry details.
+It significantly improves its consistency scores in our experiments.
+For fair comparisons, we exclude these failure cases when calculating consistency and ArtScore for StylizedGS, as they could lead to very small values.
+Stylization Quality Comparisons.
+In Table 3 and 4, denotes that failure stylization cases for StylizedGS (Zhang et al., 2025) are excluded for fair comparisons.
+Please refer to our released codes for additional visual examples.
+Table 5 and Table 6 are provided as supplementary quantitative results to Table 4.
+Table 5: Following StyleID (Chung et al., 2024), we additionally report sylization quality metrics, FID, LPIPS, LPIPS-gray, CFSD, color matching loss (HistoGAN loos), as supplementary to Table 4.
+Method Train Truck
+[…]
+StyleGaussian 34.59 0.483 0.377 0.190 0.418 28.53 0.522 0.367 0.131 0.382
+G-Style 14.80 0.471 0.364 0.165 0.185 13.65 0.512 0.353 0.087 0.176
+StylizedGS 22.90 0.707 0.648 0.216 0.396 22.65 0.779 0.713 0.084 0.417
+SGSST 24.15 0.520 0.409 0.220 0.257 19.50 0.577 0.445 0.177 0.196
+Styl3R 19.77 0.670 0.598 0.244 0.364 19.28 0.682 0.575 0.102 0.350
+Stylos (ours) 15.30 0.620 0.529 0.223 0.241 16.67 0.625 0.471 0.084 0.237
+Table 6: Following StyleID (Chung et al., 2024), we additionally report sylization quality metrics, FID, LPIPS, LPIPS-gray, CFSD, color matching loss (HistoGAN loos), as supplementary to Table 4.
+Method M60 Garden
+[…]
+StyleGaussian 30.54 0.506 0.413 0.138 0.467 25.23 0.569 0.454 0.189 0.480
+G-Style 13.93 0.498 0.395 0.092 0.208 16.17 0.500 0.364 0.103 0.179
+StylizedGS 28.33 0.815 0.728 0.102 0.443 33.73 0.876 0.827 0.074 0.570
+SGSST 23.95 0.552 0.458 0.204 0.264 20.93 0.529 0.415 0.233 0.228
+Styl3R 17.14 0.646 0.573 0.124 0.314 22.38 0.637 0.556 0.097 0.335
+Stylos (ours) 16.61 0.558 0.457 0.098 0.252 16.26 0.625 0.509 0.080 0.242
+Stylization Time Comparisons.
+We run 48 samples as a hardware warm-up phase, and calculate the average inference speeds over the following 112 samples.
+For fair comparisons, all methods are evaluated on the same hardware, using an NVIDIA GH200 GPU with identical configurations.
+The reported stylization time of the per-scene fitting required method considers per-scene training time besides the rendering time.
+We strictly follow the original authors’ training setups and inference pipelines.
+All 3DGS-based methods (Liu et al., 2024; Galerne et al., 2025; Zhang et al., 2025; Kovács et al., 2024) are evaluated on the full-resolution images, whereas single-forward methods, including Styl3R (Wang et al., 2025b) and Stylos, are evaluated at their preset input resolutions, i.e., $256\times 256$ and $448\times 448$, respectively.
+(a) The reconstructed metal bars on the gate are unclear as shown in the mid row, and the stylized metal bars at the top row are barely visible either.
+(b) The green colors fail to correspond to the intended regions in the reconstructed scene (at the mid row), and the stylized multi-view images mistakenly interpret the green area as part of the target pattern, leading the model to apply stylization to it, (at the top row).
+Figure 11: Failures caused by inaccurate scene reconstruction, such as missing geometric details or incorrect region colors.
+Top row: The input style image and the stylized multi-view results; mid row: a reference frame from the input contents and the reconstructed multi-view images; bottom row: the input contents.
+(a) The reconstructed scene is generally accurate. However, our model introduces a pure white region in the bottom-left area (at the top row), which does not appear in the original style image.
+(b) The reconstructed scene is generally accurate. However, our model hallucinates a pure white region on the left side (at the top row), which is absent from the original style image.
+Figure 12: Failures caused by generating colors that are absent from the original style images.
+Top row: The input style image and the stylized multi-view results; mid row: a reference frame from the input contents and the reconstructed multi-view images; bottom row: the input contents.
+(a) Our model produces a blurred stylized region in the upper-left area of the results (as shown in the top row).
+Figure 13: Failures caused by introducing blurry or over-smoothed style colors.
+Top row: The input style image and the stylized multi-view results; mid row: a reference frame from the input contents and the reconstructed multi-view images; bottom row: the input contents.
+A.5 Limitations
+Stylos tends to underperform on scenes dominated by high-frequency or highly cluttered structures, such as dense foliage, thin branches, or wire-like elements, where preserving fine geometric detail is particularly challenging.
+The method also struggles with style categories that introduce strong global lighting shifts or extreme color palettes, which can lead to over-saturation or loss of subtle appearance cues.
+In addition, performance decreases as the number of input views increases.
+This degradation primarily stems from limitations of the VGGT backbone: as view count grows, its underlying geometric reconstruction becomes less stable, which in turn negatively affects the quality of the stylized outputs.
+More specifically, the failure cases can be grouped into three categories: (1) Failures caused by inaccurate scene reconstruction, such as missing geometric details or incorrect region colors; (2) Failures caused by generating colors that are absent from the original style images.
+(3) Failures caused by introducing blurry or over-smoothed style colors.
+Figure 14: Style images from WikiArt (2025) used for model evaluation in all quantitative comparisons, while not used for training Stylos.
+Figure 15: Style images from WikiArt (2025) used for model evaluation in all quantitative comparisons, while not used for training Stylos.
+A.6 Additional Visual Results
+Figure 14-15 exhibit all the 50 test style images for quantitative comparisons.
+Figure 16-25 presents additional qualitative results of Stylos on diverse style–content pairs.
 Across different styles and object categories, Stylos produces visually pleasing stylizations that respect the input content geometry while transferring the target artistic appearance.
-These examples further demonstrate the versatility and robustness of our approach beyond the main results shown in the paper.
-Figure 7: Additional visual results of Stylos on diverse style–content pairs on Tanks and Temples Dataset (Knapitsch et al.
-(2017a)).
+These examples further demonstrate the versatility and robustness of our approach beyond the main results in the paper.
+Figure 16: Additional visual results of Stylos on diverse style–content pairs on Tanks and Temples Dataset (Knapitsch et al., 2017).
 Stylos consistently produces visually pleasing stylizations that preserve content geometry while transferring the target style.